4. Write the following numbers and results in standard notation, with appropriate significant figures. a. 7.85 x 103 b. 7.85 x 104 c. 5.920 x 103 d. 7.85 x10+10 e. 7.00 x 10-5 f. circumference of a circle, 2tr, where r 8.7 cm x 6.02214x 10 5.23 x 10-6 6.626x10 2.9979x 10 g.

I only need help with f. and g.
I know the rest of them. Thank you!
F and G
Answer #1

Hi,

Answer f Circumference of the circle = 2*pi*r = 2*3.14*8.7 = 54.63 = 54.6 cm, since answer can not be more precise than the least precise measurement, r = 8.7

Answer g: value of the term = 22.87218 * 103, again since the least precise value is 5.23 in the denominator, the significant value would be: 22.87*103 = 22870

Thanks, please share your feedback.